# Personality Alignment of Large Language Models

### v1

@@ Abstract
Current methods for aligning large language models (LLMs) typically aim to reflect general human values and behaviors, but they often fail to capture the unique characteristics and preferences of individual users. To address this gap, we introduce the concept of Personality Alignment. This approach tailors LLMs’ responses and decisions to match the specific preferences of individual users or closely related groups. Inspired by psychometrics, we created the Personality Alignment with Personality Inventories (PAPI) dataset, which includes data from 300,000 real subjects, each providing behavioral preferences based on the Big Five Personality Factors. This dataset allows us to quantitatively evaluate the extent to which LLMs can align with each subject’s behavioral patterns. Recognizing the challenges of personality alignments—such as limited personal data, diverse preferences, and scalability requirements—we developed an activation intervention optimization method. This method enhances LLMs’ ability to efficiently align with individual behavioral preferences using minimal data and computational resources. Remarkably, our method, PAS, achieves superior performance while requiring only 1/5 of the optimization time compared to DPO, offering practical value for personality alignment. Our work paves the way for future AI systems to make decisions and reason in truly personality ways, enhancing the relevance and meaning of AI interactions for each user and advancing human-centered artificial intelligence. The code has released in https://github.com/zhu-minjun/PAlign.
@@ 1 Introduction
The alignment with human preferences has been a focal point of both theoretical and applied research for AI systems (Shen et al., 2023; Ji et al., 2023; Hendrycks et al., 2023). However, existing alignment approaches often treat society values as a monolithic construct, lacking the granularity needed to cater to individual differences. This broad approach may overlook the nuanced preference of individual users or closely related groups, leading to a one-size-fits-all model that fails to effectively serve diverse user needs (Fernandes et al., 2023; Carranza et al., 2023; van Wynsberghe, 2021). This gap is particularly critical as personality alignment can enhance user satisfaction, trust, and engagement by making interactions more relevant and meaningful. As shown in Figure 1, we propose "Personality Alignment":
@@
As large language models (LLMs) advance (Solaiman et al., 2019; Farina & Lavazza, 2023), AI systems increasingly exhibit societal values and human-like personality traits (Ng et al., 2024; Bai et al., 2022b; Durmus et al., 2023; Zhang et al., 2024; tse Huang et al., 2024; Jang et al., 2023a). The Role-playing tasks (Li et al., 2023; Salemi et al., 2024; Gupta et al., 2024b; Yu et al., 2024) highlight this trend. However, human preferences are complex, every individual has unique preferences and needs that must be considered to effectively meet user expectations in real-world applications (Pfeifer & Bongard, 2006). For example, in customer service, some users prioritize politeness, while others prefer efficiency. Therefore, personalization is crucial for aligning AI systems with user preferences. Traditional Role-Playing Agents simulate specific professions or roles, rather than deeply aligning with individual preferences (Chen et al., 2024; Kulveit, 2018; Hubinger et al., 2019; Stray, 2020). The lack of large-scale, theoretically supported datasets for aligning AI with diverse human personalities and preferences underscores the need for further research in personality and personality alignment (Jang et al., 2023b; Lou et al., 2024; Li et al., 2024b).
There remains a practical challenge: aligning a model like Llama-2 typically requires around 3 million sentences (Touvron et al., 2023a). For personality alignment, collecting such an enormous set of behavioral data from a large number of individuals is both expensive and unrealistic. Instead, we propose a training-free personal-alignment paradigm, which provides over 300,000 unique preference samples, and a method efficiently aligning individual preference characteristics during the model’s forward computation. Specifically, we utilize the IPIP-NEO-120 questionnaire (Johnson, 2014), including 120 multiple-choice questions that quantify individual behavior patterns across five dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (De Raad, 2000; Roccas et al., 2002). This questionnaire allows subjects to provide a detailed measure of their personality and behavior within 10-20 minutes, making personality alignment both feasible and efficient (Aluja & Garcia, 2004). We collect preference data from over 300,000 subjects who complete two questionnaires: the IPIP-NEO-120 and the IPIP-NEO-300 (Goldberg et al., 1999). The IPIP-NEO-300, a superset of the IPIP-NEO-120, offers a broader range of behavioral descriptions with 300 multiple-choice questions. We manipulate the original dataset to build the Personality Alignment with Personality Inventories (PAPI) dataset, requiring LLMs to learn from each person’s IPIP-NEO-120 answers and generalize to IPIP-NEO-300 questionnaire. We then selected 300 subjects as a representative Test-set and released the remaining data as a development set, providing a novel evaluation framework and metrics for evaluating LLMs’ personality alignment fairly.
@@
To manage the behaviors of millions, the system must be scalable, ensuring quality and efficient alignment. Traditional RLHF techniques, which lack data and computational efficiency (Yao et al., 2023), are unsuitable for this task. Similarly, Prompt Engineering, though efficiently, often varies too greatly between models to be reliable for precise alignment (Zhou et al., 2023b). Inspired by activation engineering (Turner et al., 2023), we further modify the inference-time intervention (ITI)(Li et al., 2024a) by simultaneously searching for the activation value offset direction and offset distance across multiple personality dimensions. We propose a simple and effective personality alignment method: Personality Activation Search. PAS identifies the most effective attention heads in language models (LLMs) using established behavioral patterns that correlate with the five major personality traits. It then adjusts the activations in these specific directions and optimizes the magnitude of movement to achieve the best results. This approach is computationally efficient because it doesn’t rely on backpropagation to modify model parameters. Instead, PAS offers a minimally invasive intervention, resulting in greater data efficiency and more controlled personality alignment than traditional Prompt Engineering (Jiang et al., 2024).
We present a novel systematic and comprehensive personality alignment framework that enables LLMs to understand and align precisely with specific values, preferences, and behaviors required. Compared to aligning with universal values, personality alignment requires LLMs to align with individual values, preferences, and behaviors, which ensures that the LLMs can accurately match or mimic the specific individual characteristics that users need and pursue. To achieve this goal, we introduce the PAPI dataset, which covers over 300,000 subjects of different ages, genders, and periods using a multiple-choice question format to measure the alignment results. Based on that, we propose the PAS method, which aims to address the challenges of scarce individual characteristic data and high scalability in personality alignment. Experiments have shown its high alignment efficiency compared to DPO and PPO (Schulman et al., 2017; Rafailov et al., 2023) where it requires only 1/6 of the time to achieve superior performance. It even outperforms the GPT-4o model using PAS based on the Llama-3-8B model.
@@ 2 Related Work
Preferences of language models. The community of LLMs has started using personality testing tools for both qualitative and quantitative evaluations (Hoover et al., 2019; Jiang et al., 2022; Qiu et al., 2022; Xu et al., 2023; Salemi et al., 2024; Wang et al., 2024b). For example, the Machine Personality Inventory (MPI) Jiang et al. (2024), utilizes the Big-Five Personality Factors to standardize the assessment of LLMs. This approach analyzes the behaviors of LLMs across the five dimensions of the Big-Five, providing a quantitative interpretation of the values and personality traits of LLMs. Such tools are grounded in widely accepted theories, including the Big-Five and the 16 Personality Factors (Cattell & Mead, 2008), helping to contextualize LLMs’ personality traits in terms of human individual differences (Gupta et al., 2024a; Wang et al., 2023; Aher et al., 2023; Liu et al., 2024a; Li et al., 2023). These studies indicate that aligned LLMs share similar values with humans (Mehta et al., 2020; Pan & Zeng, 2023; Mao et al., 2023; Go et al., 2023; Chen et al., 2024). Jang et al. (2023b) has also mentioned the similar term "personality alignment", which considers a limited scope of preference personalization (Wang et al., 2024a; Zhuang et al., 2024; Zeng et al., 2023), focusing on only 8 different conflicting user preferences derived from combinations of three dimensions: Expertise (elementary vs. expert), Informativeness (concise vs. informative), and Style (friendly vs. unfriendly). Our work extends these theories and methods by leveraging psychological measurement to assess the value alignment gap between aligned LLMs and individual users. We aim to understand how closely aligned LLMs are with individual values through personality and personalized adjustments.
Interventions Activate. Activation interventions in neural networks trigger specific changes in internal activations (Olsson et al., 2022; Wu et al., 2024a) and have emerged as a pivotal tool in model robustness (He et al., 2019), editing (Meng et al., 2022), circuit finding (Goldowsky-Dill et al., 2023), and knowledge tracing (Geva et al., 2023). These interventions are advantageous due to their adjustable nature and minimal invasiveness (Wu et al., 2024b). Prior research has demonstrated that activation interventions can effectively identify the crucial heads and directions in a model’s internal representations (Burns et al., 2022; Li et al., 2024a; Liu et al., 2024b), which offers better data efficiency and faster optimization than traditional reinforcement learning approaches (Turner et al., 2023; Rimsky et al., 2023; Shi et al., 2024; Weng et al., 2024a). Building on these findings, our proposed method enhances scalability beyond traditional reinforcement learning by pinpointing the value directions of different individuals within the model and calculating the optimal offset distance to achieve alignment.
@@ 3 Personality Alignment Dataset Construction
We introduce dataset construction process and evaluation methods to democratise research on Personality Alignment. Detailed analysis of the dataset is provided in Appendix G.
@@
Motivation. In order to systematically evaluate the consistency between Personality-aligned LLMs and human preferences, we create a large-scale real-world Personality Alignment with Personality Inventories (PAPI) dataset, which includes 307,313 samples stored in multiple-choice questionnaires. Each sample contains a participant’s answers to the psychological questionnaires, which is shown in Figure 2. Each sample contains an IPIP-NEO-120 questionnaire and an IPIP-NEO-300 questionnaire, similar to (b). Their collection process comes from questions similar to (a), which require each subject to choose behaviors that are similar to their own. Finally, we can synthesize a personality trend analysis chart for a subject based on all the questions, as shown in (c).
Question Type. Each question is associated with one of the five major personality dimensions and correlates positively or negatively, which describes a specific behavioral pattern or preference, with five response options ranging from “Very Accurate” to “Very Inaccurate.” We expect aligned model can clearly articulate its views on these descriptions. Given the stability of individual personality tendencies, consistent responses are expected for the same personality dimension descriptions. In particular, we use each subject’s answers to the IPIP-NEO-120 questionnaire for training, and their answers to the IPIP-NEO-300 questionnaire for testing. This setup aims to measure whether the model can generalize from known behavior descriptions and personality tendencies to a broader range of behaviors, thus aligning with specific preferences.
@@
Data Collection. We have collected 307,313 samples based on the IPIP-NEO-120 and IPIP-NEO-300 from the International Personality Item Pool (IPIP) [^1] https://ipip.ori.org/index.htm. This includes approximately 60% (185,149) female and 40% (122,164) male participants, covering ages from 10 to 99 years (average age 25.19 years). The data were collected through online surveys on a website, requiring participants to provide genuine feedback on 120 & 300 behavioral questions. Participants are volunteers from various industries worldwide, including the US, UK, France, India, and China, with tests conducted between 1998 and 2019. Each participant had to actively acknowledge that careless responses would invalidate the usefulness. Each survey took between 30 to 50 minutes, was anonymous, and collected no traceable or socioeconomic status-related data. We have permission from the administrators to use IPIP items, scales, and inventories for any purpose, commercial or non-commercial.
Data Processing. To streamline the evaluation of human personality traits within a vast dataset, we apply K-Means clustering to the PAPI dataset, which contains 307,313 samples. We reduce this dataset into 300 representative clusters as Test-Set, and release the remaining data as Dev-Set. We achieve this by identifying and selecting samples closest to each cluster’s centroid, which are the most representative of their respective groups. This selection process is mathematically described as:
$\min_{x\in S_{k}}|x-\mu_{k}|_{2}$ (1)
Here, $\mu_{k}$ is the centroid of cluster $k$, and $x$ is a sample within cluster $k$ ($S_{k}$). This approach minimizes the Euclidean distance between selected samples and their cluster centroids, optimizing the representativeness of each cluster in the dataset.
[…]
@@
In this formula, $D^{Test}_{d,i}$ (the answer of IPIP-NEO-300) denotes the set of behaviors related to trait $d$ in the test section for subject $i$. We use the given template to obtain responses from both the LLM and the Person, and then calculate the absolute difference between the two. $N_{d,i}$ represents the count of these behaviors. Each $\kappa$ is an independent behavior sample from this set, with the LLM’s responses evaluated against those in $D^{Train}$ (the answer of IPIP-NEO-120). The $f(.)$ represents the scoring function for multiple-choice questions. We calculated the deviation distance between the two using the absolute value method and computed the average deviation distance for all behaviors. The range for Score is from 0 to 5, where 0 means perfect alignment with individual behavior.

### v2

@@ Abstract
Aligning large language models (LLMs) typically aim to reflect general human values and behaviors, but they often fail to capture the unique characteristics and preferences of individual users. To address this gap, we introduce the concept of Personality Alignment. This approach tailors LLMs’ responses and decisions to match the specific preferences of individual users or closely related groups. Inspired by psychometrics, we created the Personality Alignment with Personality Inventories (PAPI) dataset, which includes data from over 320,000 real subjects across multiple personality assessments - including both the Big Five Personality Factors and Dark Triad traits. This comprehensive dataset enables quantitative evaluation of LLMs’ alignment capabilities across both positive and potentially problematic personality dimensions. Recognizing the challenges of personality alignments—such as limited personal data, diverse preferences, and scalability requirements—we developed an activation intervention optimization method. This method enhances LLMs’ ability to efficiently align with individual behavioral preferences using minimal data and computational resources. Remarkably, our method, PAS, achieves superior performance while requiring only 1/5 of the optimization time compared to DPO, offering practical value for personality alignment. Our work paves the way for future AI systems to make decisions and reason in truly personality ways, enhancing the relevance and meaning of AI interactions for each user and advancing human-centered artificial intelligence. The dataset and code are released at https://github.com/zhu-minjun/PAlign.
@@ 1 Introduction
The alignment with human preferences has been a focal point of both theoretical and applied research for AI systems (Shen et al., 2023; Ji et al., 2023; Hendrycks et al., 2023). However, existing alignment approaches often treat society values as a monolithic construct, lacking the granularity needed to cater to individual differences. This broad approach may overlook the nuanced preference of individual users or closely related groups, leading to a one-size-fits-all model that fails to effectively serve diverse user needs (Fernandes et al., 2023; Carranza et al., 2023; van Wynsberghe, 2021). This gap is particularly critical as personality alignment can enhance user satisfaction, trust, and engagement by making interactions more meaningful. As shown in Figure 1, we propose "Personality Alignment":
@@
As large language models (LLMs) advance (Solaiman et al., 2019; Farina & Lavazza, 2023), AI systems increasingly exhibit societal values and human-like personality traits (Ng et al., 2024; Bai et al., 2022b; Durmus et al., 2023; Zhang et al., 2024; tse Huang et al., 2024; Jang et al., 2024). The Role-playing tasks (Li et al., 2023; Salemi et al., 2024; Gupta et al., 2024b; Yu et al., 2024) highlight this trend. However, human preferences are complex, every individual has unique preferences and needs that must be considered to effectively meet user expectations in real-world applications (Pfeifer & Bongard, 2006). For example, in customer service, some users prioritize politeness, while others prefer efficiency. Therefore, personalization is crucial for aligning AI systems with user preferences. Traditional Role-Playing Agents simulate specific professions or roles, rather than deeply aligning with individual preferences (Chen et al., 2024; Kulveit, 2018; Hubinger et al., 2019; Stray, 2020). The lack of large-scale, theoretically supported datasets for aligning AI with diverse human personalities and preferences underscores the need for further research in personality and personality alignment (Jang et al., 2023; Lou et al., 2024; Li et al., 2024b).
There remains a practical challenge: aligning a model like Llama-2 typically requires around 3 million sentences (Touvron et al., 2023a). For personality alignment, collecting such an enormous set of behavioral data from individuals is both expensive and unrealistic. Instead, we propose a training-free personal-alignment paradigm, providing over 320,000 unique preference samples across two widely-used personality assessments. Specifically, we utilize the IPIP-NEO-120 questionnaire (Johnson, 2014), which quantifies individual behavior patterns across five dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (De Raad, 2000; Roccas et al., 2002). This questionnaire enables subjects to provide detailed personality measures within 10-20 minutes (Aluja & Garcia, 2004). We collect preference data from over 300,000 subjects completing both IPIP-NEO-120 and IPIP-NEO-300 (Goldberg et al., 1999) questionnaires. Additionally, we incorporate 18,192 independent samples from the Short Dark Triad (Jones & Paulhus, 2014) assessment, which measures three negative personality traits (Machiavellianism, Narcissism, and Psychopathy) through 27 questions. We manipulate these datasets to build the Personality Alignment with Personality Inventories (PAPI) dataset. For evaluation, we selected 300 representative subjects from each assessment as independent test sets, establishing a comprehensive framework for assessing LLMs’ personality alignment capabilities across both positive and negative personality dimensions.
@@
To manage the behaviors of millions, the system must be scalable, ensuring quality and efficient alignment. Traditional RLHF techniques, which lack data and computational efficiency (Yao et al., 2023), are unsuitable for this task. Similarly, Prompt Engineering, though efficiently, often varies too greatly between models to be reliable for precise alignment (Zhou et al., 2023b). While recent activation-based approaches have shown promise in steering model behaviors (Turner et al., 2023; Li et al., 2024a), they typically focus on maximizing a single objective (e.g., safety or truthfulness). In contrast, we propose a more nuanced approach that simultaneously searches for optimal activation value offset directions and distances across multiple personality dimensions, ensuring appropriate alignment levels. This balanced activation steering is crucial for maintaining authentic personality expression while achieving desired alignment. We propose a simple and effective personality alignment method: Personality Activation Search. PAS identifies the most effective attention heads in language models (LLMs) using established behavioral patterns that correlate with the five major personality traits. It then adjusts the activations in these specific directions and optimizes the magnitude of movement to achieve the best results. This approach is computationally efficient because it doesn’t rely on backpropagation to modify model parameters. Instead, PAS offers a minimally invasive intervention, resulting in greater data efficiency and more controlled personality alignment than traditional Prompt Engineering (Jiang et al., 2024).
We present a novel systematic and comprehensive personality alignment framework that enables LLMs to understand and align precisely with specific values, preferences, and behaviors required. Compared to aligning with universal values, personality alignment requires LLMs to align with individual values, preferences, and behaviors, which ensures that the LLMs can accurately match or mimic the specific individual characteristics that users need and pursue. To achieve this goal, we introduce the PAPI dataset, which covers over 320,000 subjects of different ages, genders, and periods using multiple-choice questions to comprehensively measure both positive and negative personality traits. Based on that, we propose the PAS method, which aims to address the challenges of scarce individual characteristic data and high scalability in personality alignment. Experiments have shown its high alignment efficiency compared to DPO and PPO (Schulman et al., 2017; Rafailov et al., 2023) where it requires only 1/6 of the time to achieve superior performance. It even outperforms the GPT-4o model using PAS based on the Llama-3-8B model.
@@ 2 Related Work
Preferences of language models. The community of LLMs has started using personality testing tools for both qualitative and quantitative evaluations (Hoover et al., 2019; Jiang et al., 2022; Qiu et al., 2022; Xu et al., 2023; Salemi et al., 2024; Wang et al., 2024c). For example, the Machine Personality Inventory (MPI) Jiang et al. (2024), utilizes the Big-Five Personality Factors to standardize the assessment of LLMs. This approach analyzes the behaviors of LLMs across the five dimensions of the Big-Five, providing a quantitative interpretation of the values and personality traits of LLMs. Such tools are grounded in widely accepted theories, including the Big-Five and the 16 Personality Factors (Cattell & Mead, 2008), helping to contextualize LLMs’ personality traits in terms of human individual differences (Gupta et al., 2024a; Wang et al., 2024b; Aher et al., 2023; Liu et al., 2024a; Li et al., 2023). These studies indicate that aligned LLMs share similar values with humans (Mehta et al., 2020; Pan & Zeng, 2023; Mao et al., 2023; Go et al., 2023; Chen et al., 2024). Jang et al. (2023) has also mentioned the similar term "personality alignment", which considers a limited scope of preference personalization (Wang et al., 2024a; Zhuang et al., 2024; Zeng et al., 2023), focusing on only 8 different conflicting user preferences derived from combinations of three dimensions: Expertise (elementary vs. expert), Informativeness (concise vs. informative), and Style (friendly vs. unfriendly). Our work extends these theories and methods by leveraging psychological measurement to assess the value alignment gap between aligned LLMs and individual users. We aim to understand how closely aligned LLMs are with individual values through personality and personalized adjustments.
Interventions Activate. Activation interventions in neural networks trigger specific changes in internal activations (Olsson et al., 2022; Wu et al., 2024a) and have emerged as a pivotal tool in model robustness (He et al., 2019), editing (Meng et al., 2022), circuit finding (Goldowsky-Dill et al., 2023), and knowledge tracing (Geva et al., 2023). These interventions are advantageous due to their adjustable nature and minimal invasiveness (Wu et al., 2024b). Prior research has demonstrated that activation interventions can effectively identify the crucial heads and directions in a model’s internal representations (Burns et al., 2023a; Liu et al., 2024b), which offers better data efficiency and faster optimization than traditional reinforcement learning approaches (Turner et al., 2023; Rimsky et al., 2023; Shi et al., 2024; Weng et al., 2024a). Prior activation steering work has focused on maximizing single objectives like safety (Wang & Shu, 2024; Zheng et al., 2024) or truthfulness (Li et al., 2024a). Building on these findings, our method enhances scalability by pinpointing value directions of different individuals and calculating optimal offset distances, aiming to achieve appropriate alignment for each dimension while maintaining model capabilities.
@@ 3 Personality Alignment Dataset Construction
Motivation. To systematically evaluate the consistency between Personality-aligned LLMs and human preferences, we create the large-scale Personality Alignment with Personality Inventories (PAPI) dataset, comprising 325,505 samples across multiple-choice questionnaires. The dataset uniquely combines two complementary personality assessment systems: 307,313 samples from the Big Five personality traits (IPIP-NEO assessments) and 18,192 samples from the Dark Triad inventory. While the Big Five (IPIP) samples measure adaptive personality traits through IPIP-NEO-120 and IPIP-NEO-300 questionnaires, the Dark Triad samples assess subclinical maladaptive traits (Machiavellianism, Narcissism, and Psychopathy) through 27 targeted questions. This comprehensive approach captures both healthy personality dimensions and potentially problematic traits. As shown in Figure 2, subjects choose behaviors similar to their own (a), yielding questionnaire responses (b) that collectively form detailed personality profiles (c).
Question Type. Our dataset contains two independent questionnaire systems, each using a five-point scale from "Very Accurate" to "Very Inaccurate." The first system comprises IPIP questions associated with the Big Five personality dimensions. We use each subject’s IPIP-NEO-120 questionnaire responses to predict their IPIP-NEO-300 answers, testing the model’s ability to generalize from known personality patterns to broader behavioral descriptions. The second system consists of Dark Triad assessments, where we use 18 questions (6 each for Machiavellianism, Narcissism, and Psychopathy) to align with and predict responses to the remaining 9 questions (3 per trait). Both systems evaluate the stability of individual personality tendencies and the model’s capacity to capture and generalize personality patterns, though they operate independently on distinct trait dimensions.
@@
Data Collection. We have collected 307,313 samples based on the IPIP-NEO-120 and IPIP-NEO-300 from the International Personality Item Pool (IPIP) [^1] https://ipip.ori.org/index.htm. This includes approximately 60% (185,149) female and 40% (122,164) male participants, covering ages from 10 to 99 years (average age 25.19 years). The data were collected through online surveys on a website, requiring participants to provide genuine feedback on 120 & 300 behavioral questions. Participants are volunteers from various industries worldwide, including the US, UK, France, India, and China, with tests conducted between 1998 and 2019. Each participant had to actively acknowledge that careless responses would invalidate the usefulness. Each survey took between 30 to 50 minutes, was anonymous, and collected no traceable or socioeconomic status-related data. We have permission from the administrators to use IPIP items, scales, and inventories for any purpose, commercial or non-commercial. Additionally, we collected 18,192 Dark Triad samples [^2] https://openpsychometrics.org using the same collection pipeline and quality control procedures.
Data Processing. We apply K-Means clustering separately to the IPIP and Dark Triad datasets. From the IPIP dataset (307,313 samples), we select 300 representative clusters as the first part of our Test-Set. Similarly, we cluster the Dark Triad dataset (18,192 samples) and select 300 representative samples, forming the second part of our Test-Set. The remaining data constitutes our Dev-Set. The large-scale Dev-Set captures personality variations across different ages (from teenagers to seniors), genders, cultural backgrounds (spanning multiple countries and regions), and identity groups. This comprehensive representation is crucial for developing inclusive alignment methods that respect and reflect the full spectrum of human personality diversit. Sample selection minimizes Euclidean distance to cluster centroids: $\min_{x\in S_{k}}|x-\mu_{k}|_{2}$, where $\mu_{k}$ is the centroid of cluster $k$, and $x$ is a sample within cluster $k$ ($S_{k}$), aiming to optimize cluster representativeness by minimizing the distance between samples and their respective centroids.
Evaluation Methods. Inspired by the MPI (Jiang et al., 2024), we set up a behavioral difference score as an evaluation metric. Using a psychological measurement approach, we establish a simple scoring function S: for positively correlated questions, we assign scores from 5 to 1 from “Very Accurate” to “Very Inaccurate”. We expect the aligned language model scores to be close to a specific individual’s scores, hence we use an absolute value method for measurement. We test each trait $d$ in the test section, with each participant’s behavioral difference Aligned Score calculated as follows:
@@
In this formula, $D^{Test}_{d,i}$ denotes the set of behaviors related to trait $d$ in the test section for subject $i$. We use the given template to obtain responses from both the LLM and the Person, and then calculate the absolute difference between the two. $N_{d,i}$ represents the count of these behaviors. Each $\kappa$ is an independent behavior sample from this set, with the LLM’s responses evaluated against those in $D^{Train}$. The $f(.)$ represents the scoring function for multiple-choice questions. We calculated the deviation distance between the two using the absolute value method and computed the average deviation distance for all behaviors. The range for Score is from 0 to 4, where 0 means perfect alignment with individual behavior.